# Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation

### v1

@@ 3.3.2 $\beta$-$t^{3}$VAE
3.3.3 $\tau^{2}$ Improvement
Closely analyzing the proposed derivation of $\tau^{2}$ of the $t^{3}$VAE model, we discovered a subtle issue in its mathematical formulation. The employed $\gamma$-power divergence presents a discrepancy to the correct formula in Eq. (6). We revised the formulation and the corrected $\tau^{2}$ is (See Appendix D):
$\tau^{2}$ $=\frac{1}{1+\nu^{-1}n}\left(|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}}\approx\frac{1}{1+\nu^{-1}n}\left(\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{-\frac{2\gamma}{2+\gamma m}}.$ (9)
The corrected exact form of $\tau^{2}$ is applicable when $|\Sigma_{\phi}(x)|$ is known and when the dimension of the data is low. However, for high dimensional data, as handled in this work, we get $\gamma\approx 0$. Hence, one can use the approximation without any loss in accuracy. We note that the new form of $\tau^{2}$ leads to a similar empirical value of standard deviation compared to the previous form. Nevertheless, for correctness, in our sampling from the latent space of the $t^{3}$VAE model we use the approximation in Eq. (9).
In summary, although the $t^{3}$VAE effectively models heavy-tailed distributions through Student’s t-distributions and $\gamma$-power divergence, it does not explicitly address class imbalance in the latent space by not allocating equal volume for each class. In the next section, we introduce a class-conditional variant of the $t^{3}$VAE, designed to ensure fair and balanced generation across all classes.
@@ 4 Conditional $t^{3}$-Variational Autoencoder
We propose the Conditional $t^{3}$-Variational Autoencoder (C-$t^{3}$VAE), present its formulation, training objective, and sampling strategy. C-$t^{3}$VAE models the latent space as a mixture of Student’s t-distributions, one per class, ensuring equal latent volume allocation and promoting fairness in generation. Intra-class imbalance is further handled by the heavy-tailed nature of the Student’s t-distribution prior.
@@ 4.1 Model definition
From this joint distribution, we can derive the conditional latent prior $p(z|y)=t_{m}(z|\mu_{y},\Sigma_{y},\nu)$ and decoder distribution (See Appendix B)
$p_{\theta}(x|z,y)=t_{n}\Bigg{(}x\Bigg{|}\mu_{\theta}(z),\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\Sigma_{x},\nu+m\Bigg{)}.$
Furthermore, as in $t^{3}$VAE, we define the approximate posterior $q_{\phi}(z|x)$ as a multivariate Student’s t-distribution capturing heavy-tailed structure in the latent space :
@@ 4.2 Objective function
Harnessing Eq. (5) and the defined prior and posterior distributions of the proposed C-$t^{3}$VAE, we derive in Appendix C the following class-wise objective
@@
with $C_{1}=\left(C^{\gamma}_{\nu+n,m}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2}}\frac{\nu+n+m-2}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}$ and $C_{2}=\left(\frac{C_{\nu,m+n}^{\gamma}}{|\Sigma_{x}|^{\frac{\gamma}{2}}|\Sigma_{y}|^{\frac{2\gamma+1}{2}}}\left(1+\frac{m+n}{\nu-2}\right)^{-\gamma}\right)^{\frac{1}{1+\gamma}}$. By taking $\Sigma_{x}=\sigma^{2}I$ and $\Sigma_{y}=I$, $\mathcal{L}(\gamma,y)$ objective function simplifies to :
@@
Therefore, we express the final loss function $\mathcal{L}(\gamma)$ over the whole dataset as : $\mathcal{L}(\gamma)=\sum_{y}\mathcal{L}(\gamma,y).$ As in Eq. (3), here too we consider all $p(y_{i})$ to be equal and hence rule out their contribution to the loss function. This is done to avoid emphasizing the imbalance present in the data.
@@ 4.3 Sampling distribution
Similarly to the objective function of $t^{3}$VAE, $\mathcal{L}(\gamma,y)$ in Eq. (10) can be decomposed into a reconstruction term and regularization terms. To sample from the latent space of the C-$t^{3}$VAE, we focus on the regularization terms and define the following sampling distribution :
$p_{\nu}^{\star}(z)=\sum_{y=1}^{K}\alpha_{y}\cdot t_{m}(\mu_{y},\tau^{2}I,\nu+n),\quad\forall y,\;\;\alpha_{y}=\frac{1}{K}.$ (11)
The theoretical derivation of the variance $\tau^{2}$ leads to the form expressed in Eq. (9) (Derivation in Appendix D).
The mixture-based sampling distribution we define in Eq. (11) with equal $\alpha_{y}$ ensures a uniformly sampled synthetic data across all classes, regardless of their frequency in the original training data. As a result, C-$t^{3}$VAE equipped with this sampling distribution mitigates the common problem in generative models where head-class samples dominate due to their density in latent space. Furthermore, by modifying the mixture weights $\alpha_{y}$, one can prioritize specific classes. This makes our method flexible for targeted data augmentation or fairness-aware sampling strategies.
4.3.1 $\beta$-C-$t^{3}$VAE
@@
Overall, C-$t^{3}$VAE provides a principled, flexible, and tractable framework for fair generative modeling, particularly under class-imbalanced conditions. In the following, we study the C-$t^{3}$VAE’s performance across multiple datasets with varying imbalance degrees.
@@ 5 Experiments
In this section we outline the generative performance of the proposed C-$t^{3}$VAE model on labeled datasets compared to relevant VAE baselines [^1] We restrict our comparisons to VAE-based generative models, as Latent Diffusion Models [5], the state-of-the-art in image generation, depend on VAEs as a core component. Enhancing fairness in VAEs under class imbalance is therefore a necessary precursor to advancing more complex models.. We conduct experiments on three datasets notably SVHN-LT [11], CIFAR100-LT [24, 12] and CelebA [13] each chosen to highlight different challenges.
5.1 Evaluation procedure
All models are evaluated using Fréchet Inception Distance (FID) [25], computed against a balanced test set for each dataset. This setup measures how effectively a model overcomes training set imbalance by assessing its ability to generate high-quality samples across all classes. To evaluate robustness, we impose varying degrees of imbalance during training.
For SVHN-LT and CIFAR100-LT, we introduce class imbalance by applying an exponential decay to the number of samples per class after equalizing class sizes in the original train and test sets. The imbalance ratio $\rho$ defines the ratio between the most and least frequent classes, with class-wise sample counts $M_{y_{i}}$ given by: $M_{y_{i}}=M\cdot\rho^{-\frac{y_{i}-1}{K-1}},$ where $M$ is the original per-class sample count.
For CelebA, we compute FID per attribute, treating each attribute (eg. Mustache) and its negation (eg. no Mustache) as separate binary classes. The training set uses CelebA’s inherent imbalance, while the test set is balanced by downsampling to the smaller class size. Multi-attribute generation is not considered in this work and is left for future exploration.
5.2 Results
We present both quantitative and qualitative results of the C-$t^{3}$VAE model and the models it directly improves upon notably the VAE, C-VAE and $t^{3}$VAE models with their $\beta$ variants. This controlled comparison helps isolate the contributions of key design choices:
@@ 5.2 Results
  • $t^{3}$VAE : Student’s t-distribution latent prior and $\gamma$-power divergence objective, does not use class-conditional priors, nor allows for class conditional generation. Through a comparison with this model we assess the role of conditional modeling on Student’s t-distribution priors.
@@
In the following, we analyze the latent sampling standard deviation $\tau$ employed in $t^{3}$VAE and C-$t^{3}$VAE models, varying it to assess alignment between empirical and theoretical values. Then, we present FID comparisons across baselines with optimized hyper-parameters (The tuning of $\beta$, $\nu$ and $\tau$ hyper-parameters is reported in Appendix F), followed by per-class generative evaluation.
@@
Figure 1: FID score as a function of $\tau$ for the $t^{3}$VAE and C-$t^{3}$VAE models. Results are for the imbalance ratio $\rho=100$ for the SVHN-LT and CIFAR100-LT, and for the Mustache attribute ($\rho=25$) in the case of the CelebA dataset. Other imbalance ratios’ results paint a similar picture and are provided in Appendix F.3. The horizontal dashed lines is the FID value of the best performing VAE and C-VAE on each dataset and the vertical dashed line is the value of $\tau$ as derived in Eq. (9). We note that the used models in these figures have optimized $\beta$ and $\nu$ hyper-parameters.
5.2.1 $\tau$ parameter study
Figure 1 illustrates the sampling standard deviation’s $\tau$ impact on the FID metric. We observe that models based on the Student’s t-distribution benefit from higher standard deviation on CIFAR100-LT compared to SVHN-LT. Specifically, C-$t^{3}$VAE outperforms the C-VAE for $\tau\in[0.25;0.55]$ on CIFAR100-LT and $\tau\in[0.19;0.28]$ on SVHN-LT and for all $\tau$ values on the CelebA dataset. Moreover, it surpasses the $t^{3}$VAE models’ FID for all $\tau$ values and across all datasets.
For both Student’s t-distribution based models, the optimal FID score for the SVHN-LT occurs near the theoretically derived $\tau$ value. However, for the more complex CIFAR100-LT dataset, the optimal $\tau$ is higher than the theoretical value $\tau=0.4$. We hypothesize that the higher variance reflects the model’s need during training to accommodate the greater complexity and entropy of the dataset, which makes it harder to align the learned approximate posterior with the prior and to produce confident reconstructions. For the CelebA dataset, the standard deviation parameter $\tau$ has minimal to no impact on model performance likely due to the lower variability in the dataset’s images as they all represent faces.
5.2.2 Optimized Model Results Discussion
After optimizing the hyperparameters of the various models tested in this work, we present their generation FID scores in Table 1. We provide results for $\beta$ models after optimization and non $\beta$ models ($\beta=1$) to underscore the importance of this parameter on the resulting performance as it was not explored in the $t^{3}$VAE work [10].
@@ 5.2.2 Optimized Model Results Discussion
From Table 1, $t^{3}$VAE consistently improves over the VAE model. This underlines the generative advantage of the Student’s t-distribution prior over a Gaussian one, in addition to the reconstruction advantage highlighted by [10]. Moreover, optimizing the $\beta$ hyper-parameter improves $t^{3}$VAE’s FID over the VAE on the CIFAR100-LT and CelebA, while remaining competitive on SVHN-LT. Also, qualitative results of Figure 2 show that the $t^{3}$VAE model compared to the VAE model produces significantly sharper synthetic images on the CelebA dataset.
@@
For class-conditional models, $\beta$ optimized C-$t^{3}$VAE yields strong FID improvements across all imbalance settings, surpassing all baselines. As per Table 1, it achieves a gain of up to 4, 5 and 10 FID points over the $\beta$-$t^{3}$VAE model on the imbalanced settings of the SHVN-LT, CIFAR100-LT and CelebA, respectively. This underlines the importance of an equal per-class volume in an imbalanced dataset setting. Moreover, $\beta$-C-$t^{3}$VAE reduces the FID by up to 4 and 15 points over C-VAE on SVHN-LT and CIFAR100-LT respectively as per Table 1. For the CelebA dataset, Table 1 shows that $\beta$-C-$t^{3}$VAE achieves the best results on heavily imbalanced attributes like Mustache, demonstrating better generation for underrepresented groups. This gain in performance compared to the C-VAE goes back to the use of Student’s t-distribution latent prior and its ability to better capture intra-class long-tail distributions. Additionally, qualitative samples of conditional optimized models (Figure 3) reveal sharper facial features in C-$t^{3}$VAE compared to the C-VAE. These results establish it as the most reliable model across imbalance ratios and resolutions.
In summary, C-$t^{3}$VAE after optimization of all hyper-parameters notably $\beta$, $\nu$ and $\tau$ consistently achieves the lowest FID, validating the use of class-conditional Student’s t-distribution priors and balanced latent sampling. Improvements on CIFAR100-LT are especially pronounced, highlighting the model’s robustness to severe imbalance. However, for balanced settings, C-$t^{3}$VAE remains competitive.
5.2.3 Per-class evaluation
In this section, we evaluate the conditional models on a per-class basis. However, since FID is biased on small datasets and offers limited insight as a single scalar metric, we rely on Precision, Recall, and F1 generative metrics [^2] Precision quantifies sample quality (sharpness), Recall reflects mode coverage and F1 is their harmornic mean. Together they allows us to gain deeper insight into the models’ behavior. [26]. Our results on CelebA are shown in Figure 4, with additional results for SVHN-LT and CIFAR100-LT included in Appendix G.

### v2

@@ 3.3.2 $\beta$-$t^{3}$VAE extension
In summary, although the $t^{3}$VAE effectively models heavy-tailed distributions through Student’s $t$-distributions and $\gamma$-power divergence, it does not explicitly address class imbalance in the latent space as it does not allocate equal prior mass across class-conditioned components. In the next section, we introduce a class-conditional variant of the $t^{3}$VAE, designed to enable class-balanced generation across all classes.
@@ 4 Conditional $t^{3}$-Variational Autoencoder
In this section, we propose the Conditional $t^{3}$-Variational Autoencoder (C-$t^{3}$VAE), present its formulation, training objective, and sampling strategy. C-$t^{3}$VAE models the latent space as a mixture of Student’s $t$-distributions, one per class, inducing uniform prior mass across class-conditioned components under the sampling distribution. Intra-class variability is captured through the heavy-tailed nature of the Student’s $t$ prior.
@@ 4.1 Model definition
From this joint distribution, we can derive the conditional latent prior $p(z|y)$ and decoder distribution $p_{\theta}(x|z,y)$ (See Appendix A)
$p(z|y)=t_{m}(z|\mu_{y},\Sigma_{y},\nu),\quad p_{\theta}(x|z,y)=t_{n}\Bigg(x\Bigg|\mu_{\theta}(z),\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\Sigma_{x},\nu+m\Bigg).$
By defining class-specific priors, the model ensures that each class occupies a proportionally equal latent region, countering the frequency-aligned bias of a global prior. Furthermore, as in $t^{3}$VAE, we define the posterior $q_{\phi}(z|x)$ as a multivariate Student’s $t$-distribution capturing heavy-tailed structure in the latent space :
@@
Although the posterior is defined as $q_{\phi}(z|x)$ without explicit class conditioning, the class-specific prior and objective derived in the following section enforce class-discriminative latent encodings.
@@ 4.2 Objective function
Harnessing Eq. (5) and the defined prior and posterior distributions of the proposed C-$t^{3}$VAE, we derive in Appendix B the following class-wise objective
@@
with $C_{1}=\left(C^{\gamma}_{\nu+n,m}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2}}\frac{\nu+n+m-2}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}$ and $C_{2}=\left(\frac{C_{\nu,m+n}^{\gamma}}{|\Sigma_{x}|^{\frac{\gamma}{2}}|\Sigma_{y}|^{\frac{2\gamma+1}{2}}}\left(1+\frac{m+n}{\nu-2}\right)^{-\gamma}\right)^{\frac{1}{1+\gamma}}$.
By taking $\Sigma_{x}=\sigma^{2}I$ and $\Sigma_{y}=I$, $\mathcal{L}(\gamma,y)$ objective function simplifies to :
@@
From Eq. (9), we observe that the first term aims to minimize the input reconstruction error, ensuring data fidelity, while the second term promotes alignment between the input class mean and the inferred posterior mean. The third and fourth terms jointly regularize the posterior covariance $\Sigma_{\phi}(x)$; specifically, the trace penalty restricts the total variance to encourage compactness, whereas the determinant term acts as a barrier function to prevent the covariance matrix from collapsing to a singularity.
Finally, we express the final loss function $\mathcal{L}(\gamma)$ over the whole dataset as :
$\mathcal{L}(\gamma)=\sum_{y}\mathcal{L}(\gamma,y).$
As in Eq. (3), here too we use uniform class weighting so the loss does not scale with empirical frequency and the generative step targets all classes uniformly.
@@ 4.3 Sampling distribution
To sample from the latent space of the C-$t^{3}$VAE, we define the following sampling distribution :
$p_{\nu}^{\star}(z)=\sum_{y=1}^{K}\alpha_{y}\cdot p_{\nu,y}^{\star}(z)=\sum_{y=1}^{K}\alpha_{y}\cdot t_{m}(\mu_{y},\tau^{2}I,\nu+n),\quad\forall y,\;\;\alpha_{y}=\frac{1}{K}.$ (10)
For the the variance parameter $\tau^{2}$ in $t_{m}(\mu_{y},\tau^{2}I,\nu+n)$, we derive it theoretically. In this derivation we first derive the $\gamma$-power divergence between $p_{\nu,y}^{\star}(z)=t_{m}(\mu_{y},\tau^{2}I,\nu+n)$ and $q_{\phi}(z|x)$ for every $y$. Then, we compare the obtained result to the corresponding regularization terms in $\mathcal{L}(\gamma,y)$ Eq. (9) (See Appendix C for details). The obtained form is similar to the form expressed in Eq. (8).
The mixture-based sampling distribution we define in Eq. (10) with equal $\alpha_{y}$ is not a post-hoc balancing heuristic. Rather, it is the generative counterpart of promoting uniform prior mass across class-conditioned latent components. Sampling proportionally to empirical frequency would partially reintroduce frequency-aligned bias at generation time, attenuating the structural change imposed during training. Our framework also allows prioritization by modifying the mixture weights $\alpha_{y}$ when targeted generation is desired; however, in this work we focus on the balanced setting.
4.3.1 $\beta$-C-$t^{3}$VAE extension
@@ 5 Experiments
This section presents quantitative and qualitative results for the C-$t^{3}$VAE model and closely related baselines (VAE, C-VAE, and $t^{3}$VAE, with their $\beta$ variants). This controlled comparison isolates the contributions of our design choices:
@@
  • $t^{3}$VAE : Student’s $t$-distribution latent prior and $\gamma$-power divergence objective; it does not use class-conditional priors and does not allow class-conditional generation. Comparing to this model isolates the effect of conditional modeling with heavy-tailed priors.
We also evaluate $\beta$-VAE, $\beta$-C-VAE, and $\beta$-$t^{3}$VAE variants to assess the effect of tuning regularization on latent disentanglement and generative performance.
Besides, we note that we do not aim to compete with state-of-the-art diffusion or adversarial models in raw sample quality. Instead, we isolate the effect of latent prior geometry under class imbalance within a controlled variational framework. Comparing across fundamentally different architectural families would confound inductive bias, training stability, and objectives with the geometric contribution we study. Hence, our evaluation is restricted to VAE-family models and we do not benchmark against diffusion-based generators.
@@
In the following, we first present the evaluation metrics and protocol, then analyze the latent sampling standard deviation $\tau$ used in $t^{3}$VAE and C-$t^{3}$VAE by varying $\tau$ to assess alignment between empirical and theoretical values. We then report FID comparisons across baselines with optimized hyper-parameters (the tuning of $\beta$, $\nu$, and $\tau$ is reported in Appendix E), followed by per-class generative evaluation.
5.1 Evaluation metrics
In this section, we present the evaluation metrics used throughout this work. We report FID and generative Precision, Recall and F1 because they are the most commonly used metrics in the VAE and long-tailed generation literature, enabling direct comparison with prior work. Also, these metrics allow us to assess whether the class-conditional heavy-tailed prior in C-$t^{3}$VAE improves minority-class generation relative to baselines.
5.1.1 Fréchet Inception Distance
The Fréchet Inception Distance (FID) (Heusel et al., 2017) is a standard metric for evaluating the quality of synthetic images. It measures the similarity between the distributions of real and generated images, with a score of 0 indicating identical distributions.
To compute the FID, feature encodings of real and generated images are extracted using the InceptionV3 network (Szegedy et al., 2016), excluding the classification layer. Assuming these encodings follow a multivariate Gaussian distribution, the mean vectors $\mu_{r}$ and $\mu_{s}$ and covariance matrices $\Sigma_{r}$ and $\Sigma_{s}$ are estimated for the real and synthetic image sets, respectively. The FID is then calculated as:
$FID=\|\mu_{r}-\mu_{s}\|_{2}^{2}+\operatorname{Tr}\left(\Sigma_{r}+\Sigma_{s}-2\cdot\left(\Sigma_{r}\Sigma_{s}\right)^{\frac{1}{2}}\right).$
5.1.2 Generative Precision, Recall, and F1
While FID provides a holistic measure of distributional similarity, it does not explicitly assess the quality and diversity of generated samples. To address this, generative Precision and Recall metrics have been proposed (Kynkäänniemi et al., 2019). These metrics evaluate respectively the sharpness and mode coverage of generated samples relative to the target distribution. Their harmonic mean, the F1 score, offers a balanced assessment.
Given sets of real and generated samples $X_{r}$ and $X_{g}$, feature representations are extracted using a classifier network, yielding vectors $\delta_{r}$ and $\delta_{g}$. The complete sets of feature vectors are denoted $\Delta_{r}$ and $\Delta_{g}$, with $\Delta\in\{\Delta_{r},\Delta_{g}\}$. A binary function is defined as:
$f(\delta,\Delta)=\begin{cases}1&\text{if }\|\delta-\delta^{\prime}\|_{2}\leq\|\delta^{\prime}-\text{NN}_{k}(\delta^{\prime},\Delta)\|_{2}\text{ for at least one }\delta^{\prime}\in\Delta,\\ 0&\text{otherwise,}\end{cases}$
where $\text{NN}_{k}(\delta^{\prime},\Delta)$ denotes the $k$-th nearest neighbor of $\delta^{\prime}$ in $\Delta$ (we use $k=3$ in all experiments). The function $f(\delta_{g},\Delta_{r})$ determines whether a generated sample appears realistic, while $f(\delta_{r},\Delta_{g})$ assesses whether a real sample could be reproduced by the generator. Precision and Recall are defined as:
$\text{Precision}(\Delta_{r},\Delta_{g})$ $=\frac{1}{|\Delta_{g}|}\sum_{\delta_{g}\in\Delta_{g}}f(\delta_{g},\Delta_{r}),$
$\text{Recall}(\Delta_{r},\Delta_{g})$ $=\frac{1}{|\Delta_{r}|}\sum_{\delta_{r}\in\Delta_{r}}f(\delta_{r},\Delta_{g}).$
Finally, the F1 score is computed as:
$F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$
We note that Precision, Recall, and F1 range from 0 to 1, with 1 indicating optimal performance.
5.2 Evaluation procedure
We train all models on imbalanced datasets, where the class frequency decreases as a function of the class index, and evaluate them on balanced test sets. This protocol measures robustness by assessing a model’s ability to generate high-quality samples across all classes, regardless of their frequency during training.
To evaluate performance across varying complexities, we utilize SVHN (Netzer et al., 2011), CIFAR100 (Krizhevsky, 2009; Cao et al., 2019), and CelebA (Liu et al., 2015) (see Appendix D.1). SVHN serves as a tractable baseline—simple enough to ensure convergence, yet rich enough to highlight generative discrepancies. CIFAR100 provides a more challenging benchmark with high semantic diversity, particularly testing performance in low-data regimes. Finally, CelebA enables a detailed analysis of generative quality and class balance in the presence of natural attributes.
We construct Long-Tail (LT) variants of SVHN and CIFAR100 to explicitly control the training imbalance. Specifically, we induce imbalance via an exponential decay in sample counts after equalizing the initial class sizes. The imbalance ratio $\rho$ defines the disparity between the most and least frequent classes, with the sample count $M_{y_{i}}$ for class $y_{i}$ given by:
$M_{y_{i}}=M\cdot\rho^{-\frac{y_{i}-1}{K-1}},$
where $M$ denotes the original sample count per class and $y_{i}\in\{1,\dots,K\}$ represents the class index.
For CelebA, we exploit the inherent imbalance of facial attributes. We select Mustache and Young to represent strong ($\rho\approx 25$) and moderate ($\rho\approx 3.5$) imbalance, respectively, alongside the balanced Male and Smiling attributes. We deliberately refrain from defining classes via attribute composition; such an approach would lead to a combinatorial expansion of the label space and result in prohibitive sample sparsity per class.
We employ the FID (Heusel et al., 2017) as the primary global metric to quantify the distribution shift between generated and real samples. However, as FID can be biased in low-sample regimes, we complement it with Precision, Recall, and F1-score to provide a granular, per-class evaluation of the proposed model. We detail our experimental settings in Appendix D.
5.3 $\tau$ parameter study
Figure 1 illustrates the impact of the sampling standard deviation $\tau$ on FID. Models based on the Student’s $t$-distribution benefit from higher standard deviation on CIFAR100-LT compared to SVHN-LT. Notably, C-$t^{3}$VAE outperforms C-VAE for $\tau\in[0.25;0.55]$ on CIFAR100-LT, $\tau\in[0.19;0.28]$ on SVHN-LT, and for all $\tau$ values on CelebA. Additionally, it surpasses the $t^{3}$VAE FID for all $\tau$ values and across all datasets, underlining the importance of equal per-class prior mass.
@@ 5.3 $\tau$ parameter study
Figure 1: FID score as a function of $\tau$ for the $t^{3}$VAE and C-$t^{3}$VAE models. Results are for the imbalance ratio $\rho=100$ for the SVHN-LT and CIFAR100-LT, and the Mustache attribute ($\rho=25$) for CelebA. Other imbalance ratios’ results paint a similar picture and are provided in Appendix (E.3). The horizontal dashed lines is the FID value of the best performing VAE and C-VAE on each dataset and the vertical dashed line is the value of $\tau$ as derived in Eq. (8). We note that the used models in these figures have optimized $\beta$ and $\nu$ hyper-parameters.
Moreover, for both Student’s $t$-distribution based models, the optimal FID score for SVHN-LT occurs near the theoretically derived $\tau$ value. However, for the more complex CIFAR100-LT dataset, the optimal $\tau$ is higher than the theoretical value $\tau=0.4$. The analytically derived $\tau$ provides a principled initialization grounded in divergence geometry. Dataset-specific deviations reflect encoder–decoder capacity limits rather than theoretical inconsistency, highlighting interactions between prior heaviness and representation complexity. For CelebA, $\tau$ has minimal impact on performance, likely due to the lower variability in the dataset’s images (All faces are visually similar, reducing the impact of latent scale). Developing tighter theoretical characterizations of $\tau$ that account for dataset complexity and model capacity remains an interesting direction for future work.
5.4 Optimized Model Results Discussion
After optimizing the hyperparameters of the various models tested in this work, we present their generation FID scores in Table 1. We provide results for optimized $\beta$ models and non-$\beta$ models ($\beta=1$) to underscore the importance of this parameter, which was not explored in the original $t^{3}$VAE work (Kim et al., 2024).
@@ 5.4 Optimized Model Results Discussion
From Table 1, $t^{3}$VAE improves over VAE, highlighting the generative advantage of the Student’s $t$-distribution prior over a Gaussian one, in addition to the reconstruction advantage noted in (Kim et al., 2024). Optimizing $\beta$ improves $t^{3}$VAE FID over VAE on CIFAR100-LT and CelebA, while remaining competitive on SVHN-LT. Qualitative results in Figure 2 show that $t^{3}$VAE produces sharper synthetic images on CelebA than VAE.
@@
For class-conditional models, optimized C-$t^{3}$VAE yields strong FID improvements across imbalanced settings on SVHN-LT and CIFAR100-LT, and on heavily imbalanced CelebA attributes. As shown in Table 1, it achieves gains of up to 4, 5, and 10 FID points over $\beta$-$t^{3}$VAE on imbalanced settings of SVHN-LT, CIFAR100-LT, and CelebA, respectively. This supports the role of uniform prior mass allocation in high-imbalance regimes. Moreover, $\beta$-C-$t^{3}$VAE reduces FID by up to 4 and 15 points over C-VAE on SVHN-LT and CIFAR100-LT, respectively. For CelebA, $\beta$-C-$t^{3}$VAE achieves the best results on heavily imbalanced attributes like Mustache, indicating improved generation for underrepresented classes. The gain over C-VAE follows from the Student’s $t$-distribution latent prior and its ability to better capture intra-class long-tail structure. Qualitative samples of conditional optimized models (Figure 3) show sharper facial features for C-$t^{3}$VAE than C-VAE, notably on the imbalanced Mustache attribute. Overall, C-$t^{3}$VAE exhibits the most consistent performance in high-imbalance regimes within the VAE family.
@@
5.5 Per-class evaluation
In this section, we evaluate the conditional models on a per-class basis. Since FID can be biased on small datasets and offers limited insight as a single scalar metric, we rely on Precision, Recall, and F1 metrics (Kynkäänniemi et al., 2019). Our results on CelebA are shown in Figure 4, with additional results for SVHN-LT and CIFAR100-LT included in Appendix F.